ProbabilityMedium
Question
The probability that an event A happens in one trial of an experiment is 0.4. Three independent trials of the experments are performed. The probability that the event A happens at lest ones is
Options
A.0.936
B.0.784
C.0.904
D.None of the above
Solution
Givent that, P(A) = 0.4, P
= 0.6
P (the event A happents at least one )
= 1 - P(none of the event happens)
= 1 - (0.6)(0.6)(0.6) = 1 - 0.216 = 0.784
= 0.6 P (the event A happents at least one )
= 1 - P(none of the event happens)
= 1 - (0.6)(0.6)(0.6) = 1 - 0.216 = 0.784
Create a free account to view solution
View Solution FreeMore Probability Questions
A random variable X has the probability distribution: For the events E = {X is a prime number} and F = {X...In a series of 3 independent trials the probability of exactly 2 success is 12 times as large as the probability of 3 su...India plays 2, 3 and 5 matches against Pakistan, Sri Lanka and Australia respectively. Probability that India win a matc...A digit is chosen out of digits 1, 2, 3, 4 and 5. A second digit is chosen from the remaining four digits, then the prob...For any two events A & B in a sample space :...